OpenStudy (anonymous):

In using induction to prove that the sum of the first n terms where an = (4n − 3) is n(4n − 3)/2, what would be the mathematical expression represented by Sk + ak + 1? k(4k − 2)/2+ 4k − 3 k(4k − 2)/2+ 4k − 4 k(4k − 2)/2+ 4k + 1 (4k − 2)/2 + 4k − 3

OpenStudy (loser66):

to me, it is not the correct notation. How to prove? We have to use induction to prove \(\sum_{n}^{\infty}4n-3= \dfrac{n(4n-3)}{2}\) at n=1 , the left hand side is 4*1-3=1 while the right hand side is \(\dfrac{1*(4*1-3)}{2}=\dfrac{1}{2}\) they are not equal.!!!